OpenStudy (anonymous):

cosine of theta to the second power divided by sine of theta to the second power. + csc θ sin θ

OpenStudy (anonymous):

csc=1/sin

OpenStudy (anonymous):

csc^2θ sec^2θ 1 tan^2θ

OpenStudy (agent0smith):

Hard to read... write it out using parentheses.

OpenStudy (agent0smith):

\[\Large\frac{ \cos^2 \theta }{ \sin^2 \theta }+ \csc \theta \sin \theta\]